Please help.I can't solve this problem.
A right triangle is a triangle with one angle measuring 90°. In a right triangle, the sides are related by Pythagorean Theorem,cˆ 2=aˆ2+bˆ2, where c is the hypotenuse (the side opposite the 90° angle). Find the hypotenuse when the other 2 sides’ measurements are 6 feet and 8 feet.
Answer:
Show work in this space .

Answer by venugopalramana(3286)

Please help.I can't solve this problem.
A right triangle is a triangle with one angle measuring 90°. In a right triangle, the sides are related by Pythagorean Theorem,cˆ 2=aˆ2+bˆ2, where c is the hypotenuse (the side opposite the 90° angle). Find the hypotenuse when the other 2 sides’ measurements are 6 feet and 8 feet.
Answer:
Show work in this space .
C^2=6^2+8^2=36+64=100=10*10
C=10 FEET